# FOCUS on Contamination: A Geospatial Deep Learning Framework with a Noise-Aware Loss for Surface Water PFAS Prediction

### v2

@@ Introduction
Various modeling approaches have been explored to predict contaminant distribution. Hydrological models such as SWAT (SWAT n.d.) and MODFLOW (Torcoletti 2012) simulate pollutant transport using environmental parameters like flow rates, land use, and weather conditions. While effective, these models require extensive data and are computationally expensive at large scales (Zhi et al. 2024). Kriging, a geostatistical interpolation technique, has been widely used for mapping pollutants, including in soil (Largueche 2006) and groundwater (Singh and Verma 2019). Though useful in data-sparse environments, Kriging relies on simplifying assumptions such as spatial continuity and stationarity (GISGeography 2017) that may not hold for complex contaminants like PFAS, whose transport dynamics and uncertainties are more intricate (see Section 4). Complementing these efforts, Salvatore et al. (Salvatore et al. 2022) proposed a “presumptive contamination” model that estimates PFAS presence based solely on proximity to known PFAS-using or -producing facilities. We aim to contribute to monitoring and modeling efforts by providing scalable, data-driven PFAS predictions.
@@
Recent advances in AI and geospatial modeling have opened new avenues for tackling the scalability challenges of PFAS detection. For example, DeLuca et al. (DeLuca et al. 2023) used random forests to predict PFAS contamination in fish tissue in the Columbia River Basin, leveraging geospatial data from industrial and military facilities. On a national scale, the USGS applied XGBoost to forecast PFAS occurrence in groundwater (Tokranov et al. 2023) using PFAS sources, groundwater recharge, etc. Such approaches require extensive manual feature engineering, such as aggregating data spatially before prediction is possible (Fig. 2). In addition to being resource-intensive (see Section 4), this preprocessing loses spatial context.
@@
Recent advances in geospatial deep learning have introduced foundation models trained on large-scale satellite datasets for improved spatial prediction. Models like Prithvi (Blumenfeld 2023) and SatMAE (Cong et al. 2023) use masked autoencoders (MAE) to learn generalizable spatial representations in a semi-supervised manner, leveraging both labeled and unlabeled data. Unlike traditional ML approaches that require extensive feature engineering (e.g., Fig. 2), these models process raster (image) data directly, preserving spatial dependencies. However, many of these approaches assume that key information is directly observable in satellite imagery, which is not the case for PFAS pollution (see Section 4). We therefore adapt Prithvi to use preprocessed, proxy features, such as land cover.
@@
In environmental data analysis, limited ground truth often necessitates pseudo- or augmented labels, which can introduce noise and lead to overfitting (Safonova et al. 2023). Various strategies have been proposed to mitigate these issues, including label smoothing and robust loss functions (e.g., Huber loss (Gokcesu and Gokcesu 2021), Generalized Cross-Entropy (Zhang and Sabuncu 2018), bootstrapping loss (Reed et al. 2015)), and focal loss (Lin et al. 2018)). Among these, focal loss is particularly effective for addressing class imbalance by down-weighting well-classified examples and focusing the model’s learning on hard, uncertain samples. Weakly supervised methods like the FESTA loss (Hua et al. 2022) further aim to capture spatial and feature relationships. FESTA is specifically designed for segmentation of remote sensing images using sparse annotations assuming spatial continuity, which may not hold for PFAS (see Section 4). Furthermore, our incomplete understanding of PFAS-specific hydrological processes hinders the application of both transfer and active learning approaches, making it difficult to estimate uncertainty and find reasonable proxies (Guelfo et al. 2021; Russo et al. 2020). These challenges underscore the need for methods that integrate domain-specific knowledge directly into the training process.
@@
To contextualize our data preparation pipeline, Figure 3 illustrates how ground truth samples are integrated with geospatial data to create the final dataset.
@@ Results
We first compared FOCUS against six alternative approaches for PFAS contamination prediction as shown in Figure 6. Overall, FOCUS achieves better performance across all the years and most metrics, aside from accuracy in 2019 and 2022, and precision in 2022. The Landsat-based method yielded moderate results, suggesting that geospatial data products are more effective for capturing PFAS contamination patterns than relying solely on raw satellite data. The random forest approach, despite achieving acceptable accuracy, suffered from lower F-score, precision, and recall, likely due to the limited spatial context from per-point feature aggregation. Similarly, although Kriging provided decent accuracy and precision, it struggled with recall, indicating difficulties in detecting all instances of contamination, and pollutant transport simulation demonstrated weak overall performance. The FESTA loss approach showed suboptimal performance in our experiments. This may be attributed to our dataset’s exclusive use of point data and the complex dynamics of PFAS contamination, which may challenge the spatial continuity assumptions of the FESTA loss. While rule-based masks, designed to approximate contamination using domain-specific heuristics only, can provide a reasonable estimate (and prove useful for computing noise masks, see Section FOCUS Noise Masks Improve Performance), they fall short in accuracy compared to the data-driven approach. This finding underscores the importance of incorporating domain expertise and AI for more reliable PFAS prediction.
@@
We next conducted two complementary validation efforts of predicted contamination maps of new regions – beyond regions surrounding ground truth sample points in the train and test sets – using field-collected data (see Table 1). First, our research team collected surface water samples from multiple local sites in Michigan. To reiterate, these represent previously unsampled regions and provide an independent means of assessing model generalization beyond our test set. Notably, as shown in Figure 7, the model correctly predicted 8 collected points in this evaluation. Technical details on sample analysis are available in the supplement.
@@ D Pollutant Transport Simulation
After the simulation converges, we produce a final pollutant concentration raster, thresholded by the median value of concentrations across all patches for the year to generate a binary contamination map, as illustrated in Figure 8. The binary outputs, with values of 0 and 1 representing low and high contamination respectively, visualizes the simulation’s predictions of contamination distribution. These outputs are compared against test set patches containing actual observed PFAS presence in surface water to evaluate how accurately the simulation captures observed contamination patterns. Although this standalone approach remains computationally non-trivial, it is more tractable for batch processing of multiple patches than running a full SWAT project repeatedly. Consequently, it provides a practical baseline for pollutant transport modeling within our broader framework, allowing us to compare simulated outputs to both real-world data and other modeling approaches.
@@ G Dataset
In cases where two or more patches overlapped, as illustrated in Figure 9, all such patches were assigned to the same set (either training or testing). This ensured that no part of a test patch had been seen during training, preserving the integrity of the evaluation.

### v3

@@ Introduction
Various modeling approaches have been explored to predict contaminant distribution. Hydrological models such as SWAT (SWAT n.d.) and MODFLOW (Torcoletti 2012) simulate pollutant transport using environmental parameters like flow rates, land use, and weather conditions. While effective, these models require extensive data and are computationally expensive at large scales (Zhi et al. 2024). Kriging, a geostatistical interpolation technique, has been widely used for mapping pollutants, including in soil (Largueche 2006) and groundwater (Singh and Verma 2019). Though useful in data-sparse environments, Kriging relies on simplifying assumptions such as spatial continuity and stationarity (GISGeography 2017) that may not hold for complex contaminants like PFAS, whose transport dynamics and uncertainties are more intricate (see section Results). Complementing these efforts, Salvatore et al. (Salvatore et al. 2022) proposed a “presumptive contamination” model that estimates PFAS presence based solely on proximity to known PFAS-using or -producing facilities. We aim to contribute to monitoring and modeling efforts by providing scalable, data-driven PFAS predictions.
@@
Recent advances in AI and geospatial modeling have opened new avenues for tackling the scalability challenges of PFAS detection. For example, DeLuca et al. (DeLuca et al. 2023) used random forests to predict PFAS contamination in fish tissue in the Columbia River Basin, leveraging geospatial data from industrial and military facilities. On a national scale, the USGS applied XGBoost to forecast PFAS occurrence in groundwater (Tokranov et al. 2023) using PFAS sources, groundwater recharge, etc. Such approaches require extensive manual feature engineering, such as aggregating data spatially before prediction is possible (Fig. 2). In addition to being resource-intensive (see section Results), this preprocessing loses spatial context.
@@
Recent advances in geospatial deep learning have introduced foundation models trained on large-scale satellite datasets for improved spatial prediction. Models like Prithvi (Blumenfeld 2023) and SatMAE (Cong et al. 2023) use masked autoencoders (MAE) to learn generalizable spatial representations in a semi-supervised manner, leveraging both labeled and unlabeled data. Unlike traditional ML approaches that require extensive feature engineering (e.g., Fig. 2), these models process raster (image) data directly, preserving spatial dependencies. However, many of these approaches assume that key information is directly observable in satellite imagery, which is not the case for PFAS pollution (see section Results). We therefore adapt Prithvi to use preprocessed, proxy features, such as land cover.
@@
In environmental data analysis, limited ground truth often necessitates pseudo- or augmented labels, which can introduce noise and lead to overfitting (Safonova et al. 2023). Various strategies have been proposed to mitigate these issues, including label smoothing and robust loss functions (e.g., Huber loss (Gokcesu and Gokcesu 2021), Generalized Cross-Entropy (Zhang and Sabuncu 2018), bootstrapping loss (Reed et al. 2015)), and focal loss (Lin et al. 2018)). Among these, focal loss is particularly effective for addressing class imbalance by down-weighting well-classified examples and focusing the model’s learning on hard, uncertain samples. Weakly supervised methods like the FESTA loss (Hua et al. 2022) further aim to capture spatial and feature relationships. FESTA is specifically designed for segmentation of remote sensing images using sparse annotations assuming spatial continuity, which may not hold for PFAS (see section Results). Furthermore, our incomplete understanding of PFAS-specific hydrological processes hinders the application of both transfer and active learning approaches, making it difficult to estimate uncertainty and find reasonable proxies (Guelfo et al. 2021; Russo et al. 2020). These challenges underscore the need for methods that integrate domain-specific knowledge directly into the training process.
@@
To contextualize our data preparation pipeline, Fig. 3 illustrates how ground truth samples are integrated with geospatial data to create the final dataset.
@@ Results
We first compared FOCUS against six alternative approaches for PFAS contamination prediction as shown in Fig. 6. Overall, FOCUS achieves better performance across all the years and most metrics, aside from accuracy in 2019 and 2022, and precision in 2022. The Landsat-based method yielded moderate results, suggesting that geospatial data products are more effective for capturing PFAS contamination patterns than relying solely on raw satellite data. The random forest approach, despite achieving acceptable accuracy, suffered from lower F-score, precision, and recall, likely due to the limited spatial context from per-point feature aggregation. Similarly, although Kriging provided decent accuracy and precision, it struggled with recall, indicating difficulties in detecting all instances of contamination, and pollutant transport simulation demonstrated weak overall performance. The FESTA loss approach showed suboptimal performance in our experiments. This may be attributed to our dataset’s exclusive use of point data and the complex dynamics of PFAS contamination, which may challenge the spatial continuity assumptions of the FESTA loss. While rule-based masks, designed to approximate contamination using domain-specific heuristics only, can provide a reasonable estimate (and prove useful for computing noise masks, see section Methodology), they fall short in accuracy compared to the data-driven approach. This finding underscores the importance of incorporating domain expertise and AI for more reliable PFAS prediction.
@@
We next conducted two complementary validation efforts of predicted contamination maps of new regions – beyond regions surrounding ground truth sample points in the train and test sets – using field-collected data (see Table 1). First, our research team collected surface water samples from multiple local sites in Michigan. To reiterate, these represent previously unsampled regions and provide an independent means of assessing model generalization beyond our test set. Notably, as shown in Fig. 7, the model correctly predicted 8 collected points in this evaluation. Technical details on sample analysis are available in the supplement.
@@ D Pollutant Transport Simulation
After the simulation converges, we produce a final pollutant concentration raster, thresholded by the median value of concentrations across all patches for the year to generate a binary contamination map, as illustrated in Fig. 8. The binary outputs, with values of 0 and 1 representing low and high contamination respectively, visualizes the simulation’s predictions of contamination distribution. These outputs are compared against test set patches containing actual observed PFAS presence in surface water to evaluate how accurately the simulation captures observed contamination patterns. Although this standalone approach remains computationally non-trivial, it is more tractable for batch processing of multiple patches than running a full SWAT project repeatedly. Consequently, it provides a practical baseline for pollutant transport modeling within our broader framework, allowing us to compare simulated outputs to both real-world data and other modeling approaches.
@@ G Dataset
In cases where two or more patches overlapped, as illustrated in Fig. 9, all such patches were assigned to the same set (either training or testing). This ensured that no part of a test patch had been seen during training, preserving the integrity of the evaluation.